A glass bottle full of mercury weighs 275 g. On being heated through 40oC, 1.56 g of the mercury was expelled. calculate the mass of the bottle. (Cubic expansivity of mercury, γr, is 1.8 x 10-4K-1, and linear expansivity of glass, α, is 8 x 10-6K-1)
1
Expert's answer
2020-06-15T10:24:52-0400
"m_g+m_m+m_e=275\text{ g}.\\
m_g=275-m_m-m_e=\\=275-m_m-1.56=(273.44-m_m)\text{ g},"

where "m_m" - mass of mercury left in the bottle.